OpenStudy (anonymous):

WILL FAND AND MEDAL The regression equation y = 0.5x + 1.3 approximates the number of hours it takes to arrange a dinner event, y, given the number of people attending, x. Which statement is true? a) For every extra hour taken to make arrangements for an event, the attendance increases by 0.5 people. b) For every extra hour taken to make arrangements for an event, the attendance increases by 1.3 people. c) For every extra person attending an event, the time it takes to make arrangements increases by 0.5 hour. d) For every extra person attending an event, the time it takes to make

OpenStudy (kirbykirby):

It would be C). If you have one person attending the event: you have y = 0.5(1)+1.3 = 1.8 For 2 people: y=0.5(2)+1.3 = 2.3 3 people: y = 0.5(3)+1.3 = 2.8 Notice how the number of hours increase by 0.5 every time. This in in fact because of the slope is 0.5